OpenStudy (anonymous):

y+2=5x 4x-y=-2 Are the graphs of the given equations parallel?

OpenStudy (anonymous):

no because they don't have same slope y+2=5X=2=5x-y which is 5 is slope 4x-y=-2 which is 4 is slope they are not parallel because they don't have same slope

OpenStudy (anonymous):

y=5x-2 y=4x+2 it's clear that they don't have the same slope, so they are not parallel.

OpenStudy (saifoo.khan):

yes, they arent..

OpenStudy (anonymous):

i agree